# Generalization under Byzantine & Poisoning Attacks: Tight Stability Bounds in Robust Distributed Learning

### v1

@@ D.3 CWTM aggregation rule properties
Let denote $\{{\theta_{t}\}}_{t\in\{0,\ldots,T-1\}}$ and ${\{\theta^{\prime}_{t}\}}_{t\in\{0,\ldots,T-1\}}$ the optimization trajectories resulting from two neighboring datasets $\mathcal{S},\mathcal{S}^{\prime}$, for $T$ iterations of $\mathrm{GD}$ with aggregation rule $\mathrm{SMEA}$ and learning rate $\gamma$. For $t\in\{0,\ldots,T-1\}$, let denote,
@@
Let $i\notin\mathcal{H}$, the misbehaving vectors are gradients computed on the true loss function. This enables us to exploit their regularities if $i\in S^{*}_{t}\cap S^{\prime*}_{t}$ or their boundedness if $i\in S^{*}_{t}\setminus S^{\prime*}_{t}\cup S^{\prime*}_{t}\setminus S^{*}_{t}$. This is to be contrasted with the previously studied Byzantine failure attacks—where an intermediate comparison was introduced—as we adopt a more direct bounding technique when subject to poisoning attacks. Using these relations, we derive the following bound on $\delta_{t+1}=\|\theta_{t+1}-\theta^{\prime}_{t+1}\|_{2}$:
$\delta_{t+1}\leq\quad$ $\|G^{\mathrm{GD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{GD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+\frac{2C\gamma_{t}}{(n-f)m}$
$\quad+\frac{\gamma_{t}}{n-f}\|\sum_{i\in S^{*}_{t}\setminus S^{\prime*}_{t}}\nabla\widehat{R}_{i}(\theta_{t})-\sum_{i\in S^{\prime*}_{t}\setminus S^{*}_{t}}\nabla\widehat{R}_{i}(\theta_{t}^{\prime})\|_{2}$
[…]
@@
since $|S^{*}_{t}\cap S^{\prime*}_{t}|\leq n-f$. The precise value of the expansivity coefficient depends on the regularity assumption on the loss function. We then proceed the proof, as in Section D.2.1, depending on the regularities of the underlying loss function:
@@
We now consider the $\mathrm{SGD}$ case, where for the sake of completeness we add a result for non-convex functions that was not presented in the main text.
Theorem D.11 (Stability under poisoning attacks - $\mathrm{SMEA-SGD}$).
Consider the setting described in Section 2 under poisoning attacks. Let $\mathcal{A}=\mathrm{SGD}$, with $\mathrm{SMEA}$, for $T\in\mathbb{N}^{*}$ iterations. Suppose $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ $C$-Lipschitz and $L$-smooth.
  • (i)
  • (ii)
  • (iii)
    If we assume $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ nonconvex, bounded by $\ell_{\infty}=\sup_{\theta\in\Theta,z\in\mathcal{Z}}\ell(\theta;z)$, with a monotonically non-increasing learning rate $\gamma_{t}\leq\frac{c}{Lt}$, $t\in\{0,\ldots,T-1\}$, $c>0$, we have for any neighboring datasets $\mathcal{S},\mathcal{S^{\prime}}$:
    $\sup_{z\in\mathcal{Z}}\mathbb{E}_{\mathcal{A}}[|\ell(\mathcal{A(S)};z)-\ell(\mathcal{A(S^{\prime})};z)|]\leq 2{\left(\frac{2C^{2}}{L}\frac{f+\frac{1}{m}}{n-f}\right)}^{\frac{1}{c+1}}{\left(\frac{\ell_{\infty}T}{m}\right)}^{\frac{c}{c+1}}.$
Proof.
Let denote $\{{\theta_{t}\}}_{t\in\{0,\ldots,T-1\}}$ and ${\{\theta^{\prime}_{t}\}}_{t\in\{0,\ldots,T-1\}}$ the optimization trajectories resulting from two neighboring datasets $\mathcal{S},\mathcal{S}^{\prime}$, for $T$ iterations of $\mathrm{SGD}$ with aggregation rule $\mathrm{SMEA}$ and learning rate schedule ${\{\gamma_{t}\}}_{t\in\{0,\ldots,T-1\}}$. The proof proceeds by analyzing our robust aggregation as a classical stochastic gradient descent, albeit with a random subset selection at each iteration. Notably, the poisoned vectors are treated as true gradients, although computed on arbitrary data. For $t\in\{0,\ldots,T-1\}$, let denote,
$z_{t}^{(i)}=\begin{cases}\text{an arbitrary poisoned data point},&\text{if }i\notin\mathcal{H}\\ z^{(i,J_{t}^{(i)})},J_{t}^{(i)}\text{ sampled uniformly from }\{1,\ldots,m\}&\text{if }i\in\mathcal{H}\end{cases}$
and
with $\overline{g}_{S}=\frac{1}{|S|}\sum_{i\in S}g^{(i)}_{t}$ where $g^{(i)}_{t}=\nabla\ell(\theta_{t};z^{(i)}_{t})$ and the prime notation ${(\cdot)}^{\prime}$ denotes the corresponding quantities for $\mathcal{S}^{\prime}$. At each step $t\in\{t,\ldots,T-1\}$, the intersection and differences of the two subsets $S^{*}_{t}$ and $S^{\prime*}_{t}$ verifies the following properties:
and let denote,
$G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})=\theta_{t}-\frac{\gamma_{t}}{n-f}\sum_{i\in S^{*}_{t}\cap S^{\prime*}_{t}}g^{(i)}_{t},$
and
$G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})=\theta^{\prime}_{t}-\frac{\gamma_{t}}{n-f}\sum_{i\in S^{*}_{t}\cap S^{\prime*}_{t}}g^{(i)\prime}_{t}.$
Let $i\notin\mathcal{H}$, the misbehaving vectors are gradients computed on the true loss function. This enables us to exploit their regularities if $i\in S^{*}_{t}\cap S^{\prime*}_{t}$ or their boundedness if $i\in S^{*}_{t}\setminus S^{\prime*}_{t}\cup S^{\prime*}_{t}\setminus S^{*}_{t}$. This is to be contrasted with the previously studied Byzantine failure attacks—where an intermediate comparison was introduced—as we adopt a more direct bounding technique when subject to poisoning attacks. Using these relations, we derive the following bound on $\delta_{t+1}=\|\theta_{t+1}-\theta^{\prime}_{t+1}\|_{2}$, with probability $1-\frac{1}{m}$:
$\delta_{t+1}$ $\leq\|G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}$
$\leq\|G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+2\gamma_{t}\frac{f}{n-f}C,$
@@
$\delta_{t+1}\leq\quad$ $\|G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{SGD}\prime}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}$
[…]
where the second inequality corresponds to the worst-case scenario, namely when the index of the worker with the differing data samples lies in the intersection of the two selected subsets of workers, i.e., $a\in S^{*}_{t}\cap S^{\prime*}_{t}$. Consequently, from Lemma B.2, we obtain the following recursion:
$\mathbb{E}_{\mathcal{A}}\delta_{t+1}\leq\eta^{\mathrm{SGD}}_{G_{\gamma_{t}}}\mathbb{E}_{\mathcal{A}}\delta_{t}+2\gamma_{t}\frac{f+\frac{1}{m}}{n-f}C$
where we bound
$\eta^{\mathrm{SGD}}_{G_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}}=\frac{|S^{*}_{t}\cap S^{\prime*}_{t}|}{n-f}\eta^{\mathrm{SGD}}_{G_{\gamma_{t}}}\leq\eta^{\mathrm{SGD}}_{G_{\gamma_{t}}},$
since $|S^{*}_{t}\cap S^{\prime*}_{t}|\leq n-f$. The precise value of the expansivity coefficient depends on the regularity assumption on the loss function. We then proceed the proof, as in Section D.2.1, depending on the regularities of the underlying loss function:
  • (i)
    In the Lipschitz, smooth and convex case, with $\gamma\leq\frac{2}{L}$, we have for $t\in\{0,\ldots,T-1\}$:
    $\mathbb{E}_{\mathcal{A}}\delta_{t+1}\leq\mathbb{E}_{\mathcal{A}}\delta_{t}+2\gamma\frac{f+\frac{1}{m}}{n-f}C$
    Unrolling the recursion and using the Lipschitz continuity of the loss function leads to:
    $\mathbb{E}_{\mathcal{A}}[|\ell(\theta_{T};z)-\ell(\theta_{T}^{\prime};z)|]\leq 2\gamma C^{2}T\frac{f+\frac{1}{m}}{n-f}$
  • (ii)
    In the Lipschitz, smooth and strongly convex case, with $\gamma\leq\frac{1}{L}$, we have for $t\in\{0,\ldots,T-1\}$:
    $\mathbb{E}_{\mathcal{A}}\delta_{t+1}\leq(1-\mu\gamma)\mathbb{E}_{\mathcal{A}}\delta_{t}+2\gamma\frac{f+\frac{1}{m}}{n-f}C$
    Unrolling the recursion and using the Lipschitz continuity of the loss function yields:
    $\mathbb{E}_{\mathcal{A}}[|\ell(\theta_{T};z)-\ell(\theta_{T}^{\prime};z)|]\leq 2\gamma C^{2}\frac{f+\frac{1}{m}}{n-f}\frac{1}{\gamma\mu}\left[1-{\left(1-\gamma\mu\right)}^{T}\right]\leq\frac{2C^{2}}{\mu}\frac{f+\frac{1}{m}}{n-f}$
  • (iii)
    If the loss function is Lipschitz, smooth, and nonconvex, $c>0$, $\gamma_{t}\leq\frac{c}{Lt}$, with $t_{0}\in\{0,\ldots,T-1\}$, for all $t\in\{t_{0},\ldots,T-1\}$, we obtain:
    $\mathbb{E}_{\mathcal{A}}[\delta_{t+1}|\delta_{t_{0}}=0]\leq(1+\frac{c}{t})\mathbb{E}_{\mathcal{A}}[\delta_{t}|\delta_{t_{0}}=0]+\frac{2c}{Lt}\frac{f+\frac{1}{m}}{n-f}C$
    Building on Lemma D.4, and following the same derivation as in Section D.2.2, we unroll the recursion, minimize over $t_{0}$, and obtain:
    $\mathbb{E}_{\mathcal{A}}[|\ell(\theta_{T};z)-\ell(\theta_{T}^{\prime};z)|]\leq 2{\left(\frac{2C^{2}}{L}\frac{f+\frac{1}{m}}{n-f}\right)}^{\frac{1}{c+1}}{\left(\frac{\ell_{\infty}T}{m}\right)}^{\frac{c}{c+1}}.\qed$
@@
We can actually achieve the same theoretical bound as the non-misbehaving workers framework without the robust aggregation rule. Specifically, by reverting to the classical non-robust mean aggregation, we are able to handle terms as described in [23], since data heterogeneity does not manifest in the worst-case uniform algorithmic stability analysis. However, it is important to note that while this approach simplifies the analysis, the population risk may still remain unbounded due to the inherently unbounded optimization error that can arise from mean aggregation when facing an arbitrarily poisoned dataset.
D.4.2 Proof of Theorem 3.5
In the following, we derive sharper uniform stability upper bound with $\mathrm{CWTM}$ aggregation rule for smooth and non-convex loss functions. This improvement—particularly in its dependence on the number of local samples—is enabled by aggregation rules that preserve the smoothness of the loss function
Theorem 3.5 (Stability under poisoning attacks - $\mathrm{CWTM-SGD}$).
Consider the setting described in Section 2 under poisoning attacks. Let $\mathcal{A}=\mathrm{SGD}$, with $\mathrm{CWTM}$. Suppose $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations and that there exist $\nu>0$ constant such that $n\geq(2+\nu)f$ — that is, $f/n$ is strongly bounded away from $1/2$. Let we assume $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ bounded by $\ell_{\infty}$, nonconvex, $C$-Lipschitz and $L$-smooth. Then, with a monotonically non-increasing learning rate, $\gamma_{t}\leq\frac{\nu}{2+\nu}\frac{c}{Lt}$, $t\in\{0,\ldots,T-1\}$, $c>0$, we have for any neighboring datasets $\mathcal{S},\mathcal{S^{\prime}}$:
$\sup_{z\in\mathcal{Z}}\mathbb{E}_{\mathcal{A}}[|\ell(\mathcal{A}(S);z)-\ell(\mathcal{A}(S^{\prime});z)|]\leq 2{\left(\frac{2C^{2}\nu^{2}}{{(2+\nu)}^{2}L}\right)}^{\frac{1}{c+1}}\frac{{\big{(}T\ell_{\infty}\big{)}}^{\frac{c}{c+1}}}{m{\sqrt{n}}^{\frac{1}{c+1}}}$
Proof.
Let denote $\{{\theta_{t}\}}_{t\in\{0,\ldots,T-1\}}$ and ${\{\theta^{\prime}_{t}\}}_{t\in\{0,\ldots,T-1\}}$ the optimization trajectories resulting from two neighboring datasets $\mathcal{S},\mathcal{S}^{\prime}$, for $T$ iterations of $\mathrm{SGD}$ with aggregation rule $\mathrm{CWTM}$ and learning rate schedule ${\{\gamma_{t}\}}_{t\in\{0,\ldots,T-1\}}$. Importantly, poisoned vectors are treated as legitimate gradients, despite being computed on arbitrary data. In the following, we draw a growth recursion for the sensitivity of the parameters to the perturbation introduced. Let $a,b\in\mathcal{H}\times\{1,\ldots,m\}$ the indices of the differing samples. For $t\in\{0,\ldots,T-1\}$, we denote:
$z_{t}^{(i)}=\begin{cases}\text{an arbitrary poisoned data point},&\text{if }i\notin\mathcal{H}\\ z^{(i,J_{t}^{(i)})},J_{t}^{(i)}\text{ sampled uniformly from }\{1,\ldots,m\}&\text{if }i\in\mathcal{H}\end{cases},$
$\delta_{t}=\|\theta_{t}-\theta^{\prime}_{t}\|_{2}$, $g^{(i)}_{t}=\nabla\ell(\theta_{t},z_{t}^{(i)})$, $G^{\mathrm{CWTM}}_{\gamma_{t}}(\theta_{t})=\theta_{t}-\gamma_{t}\mathrm{CWTM}(g^{(1)}_{t},\ldots,g^{(n)}_{t})$, and the prime notation ${(\cdot)}^{\prime}$ denotes the corresponding quantities for $\mathcal{S}^{\prime}$.
The proof leverages the expansivity property of the robust aggregation update without directly comparing it to the honest stochastic gradient update. Specifically, we either directly apply the expansivity result Lemma D.8 with probability $(1-\frac{1}{m})$:
$\|G^{\mathrm{CWTM}}_{\gamma}(\theta_{t})-G^{\prime\mathrm{CWTM}}_{\gamma}(\theta^{\prime}_{t})\|_{2}\leq(1+\gamma_{t}L\min\{\frac{n}{n-2f},\sqrt{n},\sqrt{d}\})\|\theta_{t}-\theta^{\prime}_{t}\|_{2}$
or perform the following calculation with probability $\frac{1}{m}$:
$\|\mathrm{CWTM}(g^{(1)}_{t},\ldots,g^{(n)}_{t})-\mathrm{CWTM}(g^{\prime(1)}_{t},\ldots,g^{\prime(n)}_{t})\|_{2}^{2}\\ \qquad=\sum_{k=1}^{d}{\left(TM({\left[g^{(1)}_{t}\right]}_{k},\ldots,{\left[g^{(n)}_{t}\right]}_{k})-TM({\left[g^{\prime(1)}_{t}\right]}_{k},\ldots,{\left[g^{\prime(n)}_{t}\right]}_{k})\right)}^{2}\hfill\\ \qquad\leq\frac{n}{{(n-2f)}^{2}}\sum_{k=1}^{d}\|({\left[g^{(1)}_{t}\right]}_{k},\ldots,{\left[g^{(n)}_{t}\right]}_{k})-({\left[g^{\prime(1)}_{t}\right]}_{k},\ldots,{\left[g^{\prime(n)}_{t}\right]}_{k})\|_{2}^{2}\hfill\\ \qquad=\frac{n}{{(n-2f)}^{2}}\sum_{k=1}^{d}\sum_{i=1}^{n}{\left({\left[g^{(i)}_{t}\right]}_{k}-{\left[g^{\prime(i)}_{t}\right]}_{k}\right)}^{2}\hfill\\ \qquad=\frac{n}{{(n-2f)}^{2}}\sum_{i=1}^{n}\|\nabla\ell(\theta_{t};z_{t}^{(i)})-\nabla\ell(\theta^{\prime}_{t};z_{t}^{\prime(i)})\|_{2}^{2}\hfill\\ \qquad=\frac{n}{{(n-2f)}^{2}}\Bigg{(}\sum_{i=1,i\neq a}^{n}\|\nabla\ell(\theta_{t};z^{(i)})-\nabla\ell(\theta^{\prime}_{t};z^{(i)})\|_{2}^{2}\hfill\\ \qquad\qquad\qquad\qquad\qquad+\|\nabla\ell(\theta_{t};z^{(a,b)})-\nabla\ell(\theta^{\prime}_{t};z^{\prime(a,b)})\|_{2}^{2}\Bigg{)}\hfill\\ \qquad\leq\frac{n}{{(n-2f)}^{2}}\left(4C^{2}+(n-1)L^{2}\|\theta_{t}-\theta^{\prime}_{t}\|_{2}^{2}\right)\hfill$
where we used Lemma D.6 in the first inequality. Using the sub-additivity of the square root we prove that we have with probability $\frac{1}{m}$:
$\|G^{\mathrm{CWTM}}_{\gamma}(\theta_{t})-G^{\prime\mathrm{CWTM}}_{\gamma}(\theta^{\prime}_{t})\|_{2}\leq(1+\frac{\gamma_{t}L\sqrt{n(n-1)}}{n-2f})\|\theta_{t}-\theta^{\prime}_{t}\|_{2}+\frac{2\gamma_{t}C\sqrt{n}}{n-2f}.$
Consequently, with $t_{0}\in\{0,\ldots,T-1\}$, we obtain the following recursion:
$\mathbb{E}_{\mathcal{A}}[\delta_{t+1}|\delta_{t_{0}}=0]$ $\leq(1+\frac{\gamma_{t}Ln}{n-2f})\mathbb{E}_{\mathcal{A}}[\delta_{t}|\delta_{t_{0}}=0]+\frac{2\gamma_{t}C\sqrt{n}}{(n-2f)m}$
$\leq(1+\frac{\gamma_{t}L(2+\nu)}{\nu})\mathbb{E}_{\mathcal{A}}[\delta_{t}|\delta_{t_{0}}=0]+\sigma_{f,m,n}\frac{c}{Lt}$
$=(1+\frac{c}{t})\mathbb{E}_{\mathcal{A}}[\delta_{t}|\delta_{t_{0}}=0]+\sigma_{f,m,n}\frac{c}{Lt}.$
where we used that as $n\geq(2+\nu)f$, then $\frac{n}{n-2f}\leq\frac{2+\nu}{\nu}$. and defined $\sigma_{f,m,n}=\frac{2C\nu\sqrt{n}}{(2+\nu)(n-2f)m}\leq\frac{2C\nu^{2}}{{(2+\nu)}^{2}\sqrt{n}m}$. We then derive our bound similarly to the case outlined in Section D.2.2. For $0\leq t_{0}\leq\min(m,T)$, $t\in\{t_{0},\ldots,T-1\}$, $\gamma_{t}\leq\frac{\nu}{2+\nu}\frac{c}{Lt}$, $c>0$, we use $1+x\leq e^{x}$ and unroll the recursion:
$\mathbb{E}_{\mathcal{A}}[\delta_{T}|\delta_{t_{0}}=0]$ $\leq\frac{c\sigma_{f,m,n}}{L}\sum_{t=t_{0}}^{T-1}\frac{1}{t}\prod_{s=t+1}^{T-1}e^{\frac{c}{s}}\leq\frac{\sigma_{f,m,n}}{L}{\left(\frac{T}{t_{0}}\right)}^{c}$
because:
[…]
Finally, considering Lemma D.4, we obtain:
$\mathbb{E}_{\mathcal{A}}[|\ell(\theta_{T};z)-\ell(\theta_{T}^{\prime};z)|]$ $\leq\min_{0\leq t_{0}\leq\min(m,T)}\frac{t_{0}\ell_{\infty}}{m}+\frac{\sigma_{f,m,n}C}{L}{\left(\frac{T}{t_{0}}\right)}^{c}$
$\leq\min_{0\leq t_{0}\leq\min(m,T)}\frac{t_{0}\ell_{\infty}}{m}+\frac{2C^{2}\nu^{2}}{{(2+\nu)}^{2}L\sqrt{n}m}{\left(\frac{T}{t_{0}}\right)}^{c}.$
We approximately minimize the expression with respect to $t_{0}$ by balancing the two terms, i.e., setting them equal:
$0\leq\widetilde{t_{0}}={\left(\frac{m\sigma_{f,m,n}C}{L\ell_{\infty}}\right)}^{\frac{1}{c+1}}T^{\frac{c}{c+1}}\underset{\textit{for sufficiently large T}\geq\frac{m\sigma_{f,m,n}C}{L\ell_{\infty}}}{\leq}T.$
Finally, we obtain the result by directly substituting $\widetilde{t_{0}}$ into the bound:
$\mathbb{E}_{\mathcal{A}}[|\ell(\theta_{T};z)-\ell(\theta_{T}^{\prime};z)|]\leq 2{\left(\frac{\sigma_{f,m,n}C}{L}\right)}^{\frac{1}{c+1}}{\left(\frac{\ell_{\infty}T}{m}\right)}^{\frac{c}{c+1}}\\ =2{\left(\frac{2C^{2}\nu^{2}}{{(2+\nu)}^{2}L\sqrt{n}m}\right)}^{\frac{1}{c+1}}{\left(\frac{\ell_{\infty}T}{m}\right)}^{\frac{c}{c+1}}=2{\left(\frac{2C^{2}\nu^{2}}{{(2+\nu)}^{2}L}\right)}^{\frac{1}{c+1}}{\left(\ell_{\infty}T\right)}^{\frac{c}{c+1}}\frac{1}{m{\sqrt{n}}^{\frac{1}{c+1}}}\qed$
D.4.3 Comparaison of upper bounds in nonconvex settings
In this section, we present additional comparisons of upper bounds in nonconvex settings that were not included in the main text. Specifically, we compare the bounds obtained using the $\mathrm{SMEA}$ aggregation rule under poisoning attacks with those obtained using $\mathrm{CWTM}$, as well as with bounds under Byzantine failure attacks.
  • (i)
    We compare $\varepsilon^{\text{Byzantine}}$ from eq. 6 to $\varepsilon^{\text{poisoning}}_{\mathrm{SMEA}}=2{\big{(}\frac{2C^{2}}{L}\frac{f+\frac{1}{m}}{n-f}\big{)}}^{\frac{1}{c+1}}{\big{(}\frac{\ell_{\infty}T}{m}\big{)}}^{\frac{c}{c+1}}$ from Theorem D.11: $\varepsilon^{\text{Byzantine}}/\varepsilon^{\text{poisoning}}_{\mathrm{SMEA}}={\big{(}\frac{\frac{1}{(n-f)m}+\sqrt{\kappa}}{\frac{1}{(n-f)m}+\frac{f}{n-f}}\big{)}}^{\frac{1}{c+1}}$. While the dependence on $m$ remains unchanged, we demonstrate an improvement from $\sqrt{\kappa}$ to $\frac{f}{n-f}$.
  • (ii)
    We compare $\varepsilon^{\text{poisoning}}_{\mathrm{SMEA}}$ from theorem D.11 to $\varepsilon^{\text{poisoning}}_{\mathrm{CWTM}}$ from eq. 7: $\varepsilon^{\text{poisoning}}_{\mathrm{SMEA}}/\varepsilon^{\text{poisoning}}_{\mathrm{CWTM}}={\big{(}\frac{{(2+\nu)}^{2}}{\nu^{2}}\big{(}\frac{\sqrt{n}}{n-f}+fm\frac{\sqrt{n}}{n-f}\big{)}\big{)}}^{\frac{1}{c+1}}.$ We observe a more favorable dependence on $m$ in the $\mathrm{CWTM}$ bound.
D.5 Proof of Theorem 3.3 for GD
In this section, we derive uniform stability lower-bound for distributed $\mathrm{GD}$ with $\mathrm{SMEA}$ aggregation rule under poisoning attacks.

### v2

@@ C.1 Expectation bounds for honest gradients’ empirical spectral norm
Let first reason about $\mathcal{A}=\mathrm{GD}$. Let denote $\{{\theta_{t}\}}_{t\in\{0,\ldots,T-1\}}$ and ${\{\theta^{\prime}_{t}\}}_{t\in\{0,\ldots,T-1\}}$ the optimization trajectories resulting from two neighboring datasets $\mathcal{S},\mathcal{S}^{\prime}$, for $T$ iterations of $\mathrm{GD}$ with aggregation rule $\mathrm{SMEA}$ and learning rate schedule ${\{\gamma_{t}\}}_{t\in\{0,\ldots,T-1\}}$. The proof proceeds, by analyzing our robust aggregation as a classical gradient descent, albeit with a random data-dependent subset selection at each iteration. Notably, the poisoned vectors are treated as true gradients, although computed on arbitrary data. For $t\in\{0,\ldots,T-1\}$, let denote,
@@
Let $i\notin\mathcal{H}$, the misbehaving vectors are gradients computed on the true loss function. This enables us to exploit their regularities if $i\in S^{*}_{t}\cap S^{\prime*}_{t}$ or their boundedness if $i\in S^{*}_{t}\setminus S^{\prime*}_{t}\cup S^{\prime*}_{t}\setminus S^{*}_{t}$. This is to be contrasted with the previously studied Byzantine failures—where an intermediate comparison was introduced—as we adopt a more direct bounding technique when subject to data poisoning. Let $a,b\in\mathcal{H}\times\{1,\ldots,m\}$ the indices of the differing samples. Using these relations, we derive the following bound on $\delta_{t+1}=\|\theta_{t+1}-\theta^{\prime}_{t+1}\|_{2}$:
$\delta_{t+1}\leq\quad$ $\|G^{\mathrm{GD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{GD}\prime}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+\frac{\gamma_{t}}{n-f}\|\sum_{i\in S^{*}_{t}\setminus S^{\prime*}_{t}}\nabla\widehat{R}_{i}(\theta_{t})-\sum_{i\in S^{\prime*}_{t}\setminus S^{*}_{t}}\nabla\widehat{R}_{i}(\theta_{t}^{\prime})\|_{2}$
$\leq\quad$ $\|G^{\mathrm{GD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{GD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+\frac{\gamma_{t}}{(n-f)m}\|\nabla\ell(\theta_{t},z^{(a,b)})-\nabla\ell(\theta_{t}^{\prime},z_{t}^{\prime(a,b)})\|_{2}$
$\leq\quad$ $\|G^{\mathrm{GD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{GD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+2\gamma_{t}C\frac{f+\frac{1}{m}}{n-f}.$
@@
since $|S^{*}_{t}\cap S^{\prime*}_{t}|\leq n-f$. The precise value of the expansivity coefficient depends on the regularity assumption on the loss function. We then solve the recursion depending on the regularity of the underlying loss function:
@@
For $\mathcal{A}=\mathrm{SGD}$, the same recursion holds for the expected distance $\delta_{t}=\mathbb{E}_{\mathcal{A}}\|\theta_{t}-\theta^{\prime}_{t}\|_{2}$ for $t\in\{1,\ldots,T\}$. We use similar notation as above, where, for $t\in\{1,\ldots,T\}$, $G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})=\theta_{t}-\gamma_{t}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}g^{(i)}_{t}$, for each $i\in\mathcal{H}$, $g^{(i)}_{t}=\nabla\ell(\theta_{t};z^{(i,J^{(i)}_{t})}_{t})$ with $J^{(i)}_{t}$ sampled uniformly from $\{1,\ldots,m\}$, and for each $i\notin\mathcal{H}$, $g^{(i)}_{t}$ is a gradient computed on poisoned data. In fact, at each iteration, a single sample is selected uniformly at random from each worker. Consequently, with probability $1-\frac{1}{m}$:
$\delta_{t+1}$ $\leq\|G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+\frac{\gamma_{t}}{n-f}\|\sum_{i\in S^{*}_{t}\setminus S^{\prime*}_{t}}g^{(i)}_{t}-\sum_{i\in S^{\prime*}_{t}\setminus S^{*}_{t}}g^{\prime(i)}_{t}\|_{2}$
$\leq\|G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+2\gamma_{t}C\frac{f}{n-f},$
@@
$\delta_{t+1}\leq\quad$ $\|G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{SGD}\prime}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+\frac{\gamma_{t}}{n-f}\|\sum_{i\in S^{*}_{t}\setminus S^{\prime*}_{t}}g^{(i)}_{t}-\sum_{i\in S^{\prime*}_{t}\setminus S^{*}_{t}}g^{\prime(i)}_{t}\|_{2}$
$\leq\quad$ $\|G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+\frac{\gamma_{t}}{n-f}\|\nabla\ell(\theta_{t},z^{(a,b)})-\nabla\ell(\theta_{t}^{\prime},z_{t}^{\prime(a,b)})\|_{2}$
$\quad+\frac{\gamma_{t}}{n-f}\|\sum_{i\in S^{*}_{t}\setminus S^{\prime*}_{t}}g^{(i)}_{t}-\sum_{i\in S^{\prime*}_{t}\setminus S^{*}_{t}}g^{\prime(i)}_{t}\|_{2}$
$\leq\quad$ $\|G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t})-G^{\mathrm{SGD}}_{\gamma_{t}|_{S^{*}_{t}\cap S^{\prime*}_{t}}}(\theta_{t}^{\prime})\|_{2}+2\gamma_{t}C\frac{f+1}{n-f},$
where the second inequality corresponds to the worst-case scenario, namely when the index of the worker with the differing data samples lies in the intersection of the two selected subsets of workers, i.e., $a\in S^{*}_{t}\cap S^{\prime*}_{t}$. Hence, with similar derivation as in the $\mathrm{GD}$ case, we obtain the following recursion:
$\mathbb{E}_{\mathcal{A}}\delta_{t+1}\leq\eta^{\mathrm{SGD}}_{G_{\gamma_{t}}}\mathbb{E}_{\mathcal{A}}\delta_{t}+2\gamma_{t}\frac{f+\frac{1}{m}}{n-f}C.$ (15)
We finish the proof as above, the depending on the regularity of the underlying loss function. ∎
@@
We can actually achieve the same theoretical bound as the non-misbehaving workers framework without the robust aggregation rule. Specifically, by reverting to the classical non-robust mean aggregation, we are able to handle terms as described in Hardt et al. (2016), since data heterogeneity does not manifest in the worst-case uniform algorithmic stability analysis. However, it is important to note that while this approach simplifies the analysis, the population risk may still remain unbounded due to the inherently unbounded optimization error that can arise from mean aggregation when facing an arbitrarily poisoned dataset.
D.2 Proof of Theorem 3.4 for GD